A −1.00-nC point charge is at the origin, and a +4.00-nC
point charge is on the
1
Expert's answer
2020-11-10T06:56:02-0500

a)


"E(3)=(9\cdot10^{9})\frac{(4\cdot10^{-9})}{(3-2)^2}-(9\cdot10^{9})\frac{(1\cdot10^{-9})}{(3-0)^2}=35\frac{N}{C}"

b)


"E(1)=-(9\cdot10^{9})\frac{(4\cdot10^{-9})}{(1-2)^2}-(9\cdot10^{9})\frac{(1\cdot10^{-9})}{(1-0)^2}\\=-45\frac{N}{C}"

c)


"E(-2)=-(9\cdot10^{9})\frac{(4\cdot10^{-9})}{(-2-2)^2}+(9\cdot10^{9})\frac{(1\cdot10^{-9})}{(-2-0)^2}\\=0\frac{N}{C}"